# Hier-SLAM++: Neuro-Symbolic Semantic SLAM with a Hierarchically Categorical Gaussian Splatting

### v1

@@ IV-E Running time
TABLE VII: Semantic performance mIoU (%) and Parameter usage (MB) on Replica. Evaluations are conducted over a total of 102 semantic classes. Results are highlighted as first, second.
@@
mIoU (%) NIDS-SLAM [12] 82.37 82.45 84.08 76.99 85.94
@@
SemGauss-SLAM [24] 96.34 96.30 95.82 96.51 96.72
SNI-SLAM [11] 87.41 88.42 87.43 86.16 87.63
@@
  • The gray background indicates that the different evaluation methods are used, where evaluations are conducted on a subset of the total classes.
IV-F Hierarchical semantic understanding
We perform semantic understanding experiments on the synthetic dataset Replica [41] to demonstrate the comprehensive performance of our proposed method. Replica [41] is a synthetic indoor dataset that includes 102 semantic classes with high-quality semantic annotations. Similar to previous works [24, 7, 11], we evaluate the mIoU performance as the quantitative semantic results in Tab. VII, where mIoU (%) is reported by comparing the rendered semantic map with the ground truth across all original semantic classes (102 classes). To further analyze the storage requirements of semantic SLAM methods, we report the parameter storage usage in the second block of Tab. VII.
From Tab. VII, our method, along with the flat coding version and the proposed one-hot version, both demonstrate superior semantic performance compared to all existing methods. By utilizing hierarchical semantic representation, Hier-SLAM++ (one-hot) achieves a similar semantic performance with less than a $1\%$ difference in mIoU while requiring 4.5× less storage compared to the flat version. The training time shown in Tab. VI demonstrates that our hierarchical version (Hier-SLAM++-one-hot) requires only 37% of the tracking time per frame while maintaining similar frame mapping performance compared to the flat version (Hier-SLAM++-flat). The comparable mIoU performance, along with significantly reduced storage and computational cost, clearly demonstrates the effectiveness and efficiency of our proposed method. Furthermore, our binary version (Hier-SLAM++-binary) achieves state-of-the-art semantic understanding while achieves the best parameter storage efficiency, reducing storage usage to 63% of the one-hot version and requiring only 20% of the storage compared to the flat version. This reduction in storage usage highlights the efficiency of our proposed method. Because SGS-SLAM [23], SemGauss-SLAM [24] and SNI-SLAM [11] use different evaluation methods that restrict the mIoU calculation to a subset of the global semantic classes. Specifically, SGS-SLAM computes the average IoU for the semantic classes visible in each frame, and SemGauss-SLAM together with SNI-SLAM [11] evaluates performance based on 52 visible classes across all evaluation sequences, making direct comparisons unfair.
[…]
IV-G Scaling up capability
@@ IV-H Monocular performance
We present our monocular SLAM performance in Tab. VIII. For camera tracking, our method achieves competitive performance compared to state-of-the-art approaches without loop closure. Specifically, compared to the state-of-the-art Gaussian Splatting-based methods [17], our method exhibits similar tracking performance while achieving superior rendering quality. For SLAM methods with loop closure, which typically demonstrate superior camera tracking performance due to well-executed loop closure mechanisms that significantly enhance tracking accuracy and overall SLAM performance, our method still achieves superior rendering accuracy compared to Photo-SLAM [19], which incorporates loop closure. This highlights the strong rendering capabilities of our proposed method. With the assistance of the geometric prior provided by the feed-forward method, the joint optimization of both photometric loss and aligned depth loss contributes to improved rendering performance in the whole SLAM system.
@@ IV-I Ablation study
We evaluate SLAM performance under different tree designs in Tab. IX. The evaluation is conducted using the baseline method, Hier-SLAM [25], ensuring that any performance differences are solely attributable to the tree structures. In Tab. IX, each metric represents the average result from two runs on the first four sequences—Room0, Room1, Room2, and Office0—from the Replica dataset. We generate two different 5-level trees using our proposed tree generation procedure for the Replica dataset and incorporate them into the semantic SLAM system. As shown in Tab. IX, all methods demonstrate good performance, which can be attributed to the well-designed tree structures generated through both LLMs and 3D Generative Models. Additionally, the fully optimized hierarchical semantic loss contributes to the robustness of SLAM performance. Among all compared methods, tree-1 achieves the best semantic mIoU performance, the highest depth mapping accuracy, and close rendering quality (PSNR, SSIM, and LPIPS) compared with other settings. Therefore, we select tree-1 as the hierarchical tree structure leveraged within our semantic SLAM system.
TABLE IX: Performance of different trees on Replica. Best results are highlighted in bold.
@@
Hier-SLAM [25] 75.517 0.292 0.416 35.885 0.982 0.057
[…]
Hier-SLAM++${}^{\text{tree2}}$ 75.061 0.280 0.413 35.792 0.982 0.057
@@ V Conclusion
We introduced Hier-SLAM++, a neuro-symbolic semantic 3D Gaussian Splatting SLAM system that supports both RGB-D and monocular inputs, incorporating a novel hierarchical categorical representation. Specifically, we proposed a compact and generalized hierarchical representation that encodes both semantic meaning and geometric attributes (i.e., size and shape), leveraging the capabilities of LLMs and 3D generative models. To enhance global semantic understanding, we introduced an improved hierarchical semantic loss. Furthermore, we extended the system from an RGB-D-only SLAM to support monocular input, utilizing geometric priors derived from a feed-forward SLAM approach. Experimental results demonstrate that Hier-SLAM++ achieves superior or on-par performance with state-of-the-art NeRF-based and Gaussian-based SLAM methods in terms of tracking, mapping, and semantic understanding accuracy, while significantly reducing storage requirements and runtime. These advances establish Hier-SLAM++ as a robust solution for semantic 3D understanding across diverse environments.
Figure 6: Visualization of RGB image rendering performances on the Replica dataset [41]. The first row presents the image rendering results of our Hier-SLAM++. The second to fourth rows show the rendering outputs of the comparison methods. The last row displays the ground truth images for reference. We use red boxes to highlight key differences regions within the image. The visual comparison clearly demonstrates that our method achieves superior rendering quality, closely resembling the ground truth images compared to state-of-the-art approaches.
@@
I Appendix 1. Implementation Details for Tree Generation
@@
No Dublications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
@@
No Dublications: make sure to not repeat any class members, the group names are fine but not the groups’ members.

### v2

@@ IV-D Rendering Quality
IV-E Running time
The runtime results for the compared methods are provided in Tab. VI, covering both NeRF-based and Gaussian Splatting-based SLAM methods. For the reason that one compared baseline using a flat representation (Hier-SLAM++-flat) cannot run successfully on an NVIDIA RTX 4090 due to GPU memory limitations, we report the runtime separately for RTX 4090 and L40S. In the first block of Tab. VI, we observe that, compared to NeRF-based methods [59], Gaussian-based state-of-the-art methods [16] achieve significantly higher operating speeds, benefiting from the fast rendering and efficient optimization capabilities inherent to Gaussian Splatting representation. Furthermore, with the carefully refined forward and backward gaussian module introduced in our system, which serves as the core component of gaussian splatting-based SLAM systems, our method (Hier-SLAM++- w/o sem) achieves up to 2.4× faster tracking and 2.2× faster mapping compared to the best-performing method [16]. In the second block of Tab. VI, we evaluate the impact of different semantic representations on runtime performance. Our method maintains high efficiency by employing hierarchical semantic coding (Hier-SLAM++-one-hot), achieving nearly 3× faster tracking compared to the flat semantic representation (Hier-SLAM++-flat). Furthermore, our proposed binary semantic representation (Hier-SLAM++-binary) demonstrates even greater efficiency, enabling more lightweight performance. The flat version is slightly faster in mapping, with an improvement of 0.38 seconds per-frame, this is due to its cross-entropy loss calculation directly without hierarchical optimization for the proposed methods. Notably, our Hier-SLAM++ achieves a rendering speed of 2000 FPS, which increases to 3000 FPS when semantic information is not included.
@@ IV-E Running time
IV-F Hierarchical semantic understanding
[…]
From Tab. VII, our method, along with the flat coding version and the proposed one-hot version, both demonstrate superior semantic performance compared to all existing methods. By utilizing hierarchical semantic representation, Hier-SLAM++ (one-hot) achieves a similar semantic performance with less than a $1\%$ difference in mIoU while requiring 4.5× less storage compared to the flat version. The training time shown in Tab. VI demonstrates that our hierarchical version (Hier-SLAM++ (one-hot)) requires only 37% of the tracking time per frame while maintaining similar frame mapping performance compared to the flat version (Hier-SLAM++ (flat)). The comparable mIoU performance, along with significantly reduced storage and computational cost, clearly demonstrates the effectiveness and efficiency of our proposed method. Furthermore, our binary version (Hier-SLAM++ (binary)) achieves state-of-the-art semantic understanding while achieves the best parameter storage efficiency, reducing storage usage to 63% of the one-hot version and requiring only 20% of the storage compared to the flat version. This reduction in storage usage highlights the efficiency of our proposed method. Since works [11, 68, 67] report mIoU only on a subset of classes, making direct comparison unfair, we follow the same evaluation protocol as [67], which computes mIoU using only the visible semantic classes in each frame. Our method achieves an mIoU of $96.79\%$ with a storage usage of $910.50$ MB, demonstrating on-par semantic rendering performance with state-of-the-art methods [68], while maintaining efficient storage usage. Meanwhile, given that [68] benefits significantly from a large foundation model pre-trained on much larger and more diverse datasets, our results demonstrate that comparable performance can be achieved without relying on such heavy models.
[…]
TABLE VII: Semantic performance mIoU (%) and Parameter usage (MB) on Replica. Results are highlighted as first, second.
@@ IV-F Hierarchical semantic understanding
mIoU (%) total 102 classes NIDS-SLAM [12] 82.37 82.45 84.08 76.99 85.94
@@
mIoU (%) subset classes SNI-SLAM [11] 87.41 88.42 87.43 86.16 87.63
SemGauss-SLAM [68] 96.34 96.30 95.82 96.51 96.72
@@
Hier-SLAM++ (one-hot) 96.79 96.93 96.78 96.92 96.53
Hier-SLAM++ (binary) 95.26 96.21 92.62 96.34 95.85
@@
  • Ours represents our method with a hierarchical representation, evaluated on a subset of semantic classes, consistent with [67].
@@ IV-H Monocular performance
We present our monocular SLAM performance in Tab. VIII. For SLAM methods without loop closure, our approach achieves competitive results compared to state-of-the-art methods in the same setting. For methods with loop closure, where well-executed loop mechanisms typically boost tracking accuracy and overall SLAM performance, our method still delivers better rendering accuracy compared to Gaussian-based SLAM systems [19]. With the assistance of the geometric prior provided by the feed-forward model, the joint optimization of both photometric loss and aligned depth loss contributes to improved rendering performance across the entire SLAM system.
@@ IV-I Ablation study
In this section, we evaluate the SLAM performance under different hierarchical tree designs. The results are presented in Tab. IX. Specifically, the experiments are conducted using our baseline setup, which adopts a one-hot hierarchical semantic representation and a naïve decoder with a single Softmax layer, denoted as Baseline in the table. This setup ensures that any observed performance differences are solely attributable to the tree structure itself. Each metric in Tab. IX represents the average result over two runs on the first four sequences—Room0, Room1, Room2, and Office0—from the Replica dataset. We generate two distinct 5-level hierarchical trees (tree1 and tree2) using our proposed tree construction framework, which integrates both LLM-generated semantics and 3D generative priors, and evaluate them within our semantic SLAM system. For comparison, we also include our previous Hier-SLAM [25] (tree0), which relies solely on LLM-generated semantics. As shown in the table, all configurations yield strong performance, which we attribute to the quality of the tree structures produced by the joint LLM and 3D prior-based generation process. In addition, the fully optimized hierarchical semantic loss contributes to the robustness of SLAM performance across tree variants. Among the compared configurations, tree2 achieves the highest semantic mIoU, the best depth reconstruction accuracy, and comparable rendering quality (in PSNR, SSIM, and LPIPS) relative to the alternatives. Therefore, we adopt tree2 as the default hierarchical structure in our semantic SLAM framework.
TABLE IX: Performance of different trees on Replica. BEST RESULTS ARE HIGHLIGHTED AS FIRST.
@@
Baseline (tree0) 75.517 0.292 0.416 35.885 0.982 0.057
Baseline (tree1) 75.061 0.280 0.413 35.792 0.982 0.057
Baseline (tree2) 76.754 0.302 0.408 35.825 0.982 0.058
We also evaluate our proposed one-hot and binary semantic embeddings in previous experiments, including camera tracking performance in Tab. I and Tab. II, reconstruction accuracy in Tab. IV, rendering quality in Tab. V, as well as runtime and semantic understanding in Tab. VI and Tab. VII, respectively. For camera tracking, mapping, and rendering performance, both embeddings exhibit similar results. However, for runtime and semantic understanding, the two embeddings show differences. Specifically, our one-hot representation achieves better semantic understanding (higher mIoU in Tab. VII), while the binary embedding shows an 8% drop in mIoU on the full semantic class evaluation and less than 1% drop on the subset evaluation, but provides faster operation and significantly reduced storage—approximately 36% savings, as illustrated in Tab. VI and Tab. VII. As introduced in Section III, our proposed semantic embeddings provide two types of compactness: the one-hot representation effectively compresses semantic information, achieving faster training and lower memory usage, while the binary embedding further reduces storage and computation time at the cost of a small decrease in semantic accuracy (which remains minimal on subset evaluations). For tracking, mapping, and rendering, the performance difference between the two embeddings is negligible, as both integrate semantic information into the SLAM process with interpretable compactness derived from our carefully designed hierarchical tree representation.
@@ V Conclusion
We introduced Hier-SLAM++, a neuro-symbolic semantic 3D Gaussian Splatting SLAM system that supports both RGB-D and monocular inputs, incorporating a novel hierarchical categorical representation. Specifically, we proposed a compact and generalized hierarchical representation that encodes both semantic meaning and geometric attributes (i.e., size and shape), leveraging the capabilities of LLMs and 3D generative models. To enhance global semantic understanding, we introduced a novel hierarchical semantic loss. Furthermore, we extended the system from an RGB-D-only SLAM to support monocular input, utilizing geometric priors derived from a feed-forward SLAM approach. Experimental results demonstrate that Hier-SLAM++ achieves superior or on-par performance with state-of-the-art NeRF-based and Gaussian-based SLAM methods in terms of tracking, mapping, and semantic understanding accuracy, while significantly reducing storage requirements and runtime. These advances establish Hier-SLAM++ as a robust solution for semantic 3D understanding across diverse environments.
@@
I Implementation Details for Tree Generation
@@
No Duplications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
@@
No Duplications: make sure to not repeat any class members, the group names are fine but not the groups’ members.
@@
II Generated tree demonstration
We demonstrate the generated hierarchical tree for the Replica dataset [78] in Fig. 1. Based on the semantic categories in Replica, we leverage the capabilities of large language models and 3D generative models to construct a hierarchical tree. This tree represents semantic information as symbolic nodes (illustrated as colored nodes), which are used to produce compact semantic embeddings for each Gaussian primitive. These embeddings are then learned end-to-end across multiple views during SLAM operation.
Figure 1: Visualization of the generated hierarchical tree on the Replica dataset [78]. Different colors represent different tree depths. Each node corresponds to a symbolic concept, with label names shown next to the nodes.